# Analysis of static and dynamic batching algorithms for graph neural networks

### v1

@@ Appendix C Mean versus median in timing measurements
For each timing experiment ten experiments are run and either the mean or median is taken. The median timing results for the MPEU model after running one hundred thousand steps is seen in Fig. 6. The mean results for the same experiments are shown in 7. The difference in the two figures shows the importance of the number of recompilations in the gradient-update method. This is because the mean is affected by outlier measurements, such as recompilations, while the median is not. The fact that for one hundred thousand training steps, the static-64 algorithm has the best median performance but not the best mean performance is due to the number of recompilations which is highest for this algorithm.
@@ Appendix E Test metric results
The test metric (RMSE) curves for each of the ten iterations are averaged for each batch size, model and dataset combination. The resulting mean curve is shown for the two models using the QM9 data in Fig. 12. We do not see significant differences in the test performance for the static-$2^{N}$ or dynamic batching algorithm. Note that we do not expect nor do we see any noticeable difference in learning from the static-64 to static-$2^{N}$ algorithms since the difference between the methods is the padding scheme which does not affect the loss. Therefore, the static-64 curve is left out of Fig. 12 for visual clarity. The results for AFLOW can be seen in the accompanying code repository.
@@ Appendix F Static-constant batching
Appendix G Node-level target discussion
Both datasets evaluated in this paper contain graph-level targets, but not node-level targets. The batching step of each algorithm does not change for a node-level target. The gradient step should also not change significantly. Graph-level targets typically feed their node feature vectors into readout functions that sum over the nodes in the graph. For node-level targets, on the other hand, the batch loss is usually computed as the sum of the losses of the node-level targets. Therefore, both target types typically contain a sum over nodes, and we expect the gradient update steps profiling results to remain similar for node-level classification. That said, the learning curves could change for node-level targets, since the true nodes that exist in a dynamic batch are sampled from a roughly truncated Gaussian (as was seen in Figure 2), whereas the static batch has a Gaussian distribution. This could have an impact on how the model learns, since batches will contain significantly different numbers of targets on average.
Appendix H Effect on rankings of the number of training steps

### v2

@@ Appendix C Mean versus median in timing measurements
For each timing experiment ten experiments are run and either the mean or median is taken. The median timing results for the MPEU model after running one hundred thousand steps is seen in Fig. 8. The mean results for the same experiments are shown in 9. The difference in the two figures shows the importance of the number of recompilations in the gradient-update method. This is because the mean is affected by outlier measurements, such as recompilations, while the median is not. The fact that for one hundred thousand training steps, the static-64 algorithm has the best median performance but not the best mean performance is due to the number of recompilations which is highest for this algorithm.
@@ Appendix E Test metric results
The test metric (RMSE) curves for each of the ten iterations are averaged for each batch size, model and dataset combination. The resulting mean curve is shown for the MPEU and SchNet models using the QM9 data in Fig. 14. We do not see significant differences in the test performance for the static-$2^{N}$ or dynamic batching algorithm. Note that we do not expect nor do we see any noticeable difference in learning from the static-64 to static-$2^{N}$ algorithms since the difference between the methods is the padding scheme which does not affect the loss. Therefore, the static-64 curve is left out of Fig. 14 for visual clarity. The results for AFLOW can be seen in the accompanying code repository.
@@ Appendix F Static-constant batching
Appendix G Effect on rankings of the number of training steps
@@
Appendix H Student t-tests on test metrics
The student t-test was performed in a pairwise fashion between two batching algorithms upon their distribution of final test RMSE values. For the SchNet model on the QM9 dataset for batch size 64, a heatmap of the t-test values is shown in Fig. 16 and the associated p-values in Fig. 17.
Figure 16: Heatmap of pairwise student t-test values on the distribution of test RMSE values for the SchNet model, QM9 data and a batch size of 64.
Figure 17: Heatmap of p-values from the pairwise student t-test on the distribution of test RMSE values for the SchNet model, QM9 data and a batch size of 64.
Figures for other datasets and models can be found in the Software and Data section.